adult tickets cost $8.50 and kids under 12 pay only $8.00. If 54 tickets are sold for a total of $445.00, how many of the tickets were adult tickets and how many were sold to kids under 12
Answer by mananth(16949)

x $ Adults tickets
y $ Under 12 kids

1 x + 1 y = 54 .............1
8.5 x + 8 y = 445 .............2
Eliminate y
multiply (1)by -8
Multiply (2) by 1
-8 x -8 y = -432
8.5 x + 8 y = 445
Add the two equations
0.5 x = 13
/ 1
x = 26
plug value of x in (1)
1 x + 1 y = 54
26 + 1 y = 54
1 y = 54 -26
1 y = 28
y = 28
$26 Adults tickets
$28 Under 12 kids
m.ananth@hotmail.ca